is the solution of 3x + 10 = 7.   
                                       is the solution of 3x + 10 = 7.

Grade:12

3 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given equation, 3x + 10 = 7
Transpose 10 to RHS then it becomes -10.
3x = 7 - 10
3x = -3
x = -3/3
x = -1
x = -1 is the solution of the equation 3x + 10 = 7.

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The solution of the equation is 
3X + 10 = 7
3X = 7-10
3X = -3
X = -3/3
X = -1
 
Hope it helps!
Thank you
mohd samiuddin
2 years ago
Given equation is 3X+10=7
send 10 to RHS side
3X=7-10
3X=-3
send 3 to RHS side
X=-3/3
 X=-1
Therefore -1 is the solution of 3X+10=7